# On the speed of coming down from infinity for (sub)critical branching processes with pairwise interactions

### v1

@@ 3.5 Proof of Theorem 4
$\mathbb{E}\Big{[}\sup_{s\leq t}(D_{s}-D_{k,s})^{2}\Big{]}\leq C\bm{\sigma}^{2}_{{\rm int},k}t^{3},\quad\text{for}\quad t\leq T,$
@@
$\langle D_{k,\cdot}-D_{\cdot}\rangle_{t}=(\mathbf{m}_{\rm int})^{2}\bm{\sigma}^{2}_{{\rm int},k}\int_{0}^{t\wedge t_{0}}s^{4}Z_{s}^{(\infty)}\left(Z_{s}^{(\infty)}-1\right){\rm d}s,\quad\text{for}\quad t\geq 0.$
Then, for any $0\leq T<\infty$, Corollary 2 (i) implies that there exists a constant $0<C<\infty$ such that
$\mathbb{E}[\langle D_{k,\cdot}-D_{\cdot}\rangle_{t}]$ $\leq C\bm{\sigma}^{2}_{{\rm int},k}t^{3},\quad\text{for}\quad t\leq T.$
@@
$D_{t}^{(\varepsilon)}:=-\varepsilon^{-\frac{3}{2}}D_{\varepsilon t},\qquad\textrm{ for }\quad t\geq 0,$
@@
$\widehat{X}_{0}=0\qquad\text{and}\qquad\widehat{X}_{t}=\bm{\sigma}_{{\rm int}}\int_{0}^{t}u{\rm d}W_{u},\qquad\text{for}\quad t>0,$
@@
$\widehat{X}_{0}^{(k)}=0\qquad\text{and}\qquad\widehat{X}_{t}^{(k)}=\hat{\bm{\sigma}}_{{\rm int},k}\int_{0}^{t}u{\rm d}W_{u},\qquad\text{for}\quad t>0,$ (90)
where $\hat{\bm{\sigma}}^{2}_{{\rm int},k}\coloneqq c+\sum_{i=1}^{\lfloor k\rfloor}i^{2}b_{i}$. Recall that convergence in $\mathbb{D}([0,\infty),\mathbb{R})$ to a continuous limit is equivalent to uniform convergence on compact subsets of $[0,\infty)$, see for instance [19, Proposition 1.17 in Chapter VI]. Therefore, since by (${\bf H4}$), $\hat{\bm{\sigma}}^{2}_{{\rm int},k}\rightarrow\bm{\sigma}^{2}_{{\rm int}}$, as $k\rightarrow\infty$, it is not difficult to see that Lemma 20 and (90) imply the result in Proposition 3.
To prove that $(D_{k,t}^{(\varepsilon)})_{t\geq 0}$ converges weakly towards $(\widehat{X}_{t}^{(k)})_{t\geq 0}$, we use [10, Theorem 1.4 in Chapter 7.1]. Note that $(D_{k,t}^{(\varepsilon)})_{t\geq 0}$ is a well-defined locally square-integrable martingale. In particular, by (89), its predictable quadratic variation $(\langle D_{k,\cdot}^{(\varepsilon)}\rangle_{t})_{t\geq 0}$ is given by
$\langle D_{k,\cdot}^{(\varepsilon)}\rangle_{t}=\varepsilon^{-3}(\mathbf{m}_{\rm int})^{2}\hat{\bm{\sigma}}^{2}_{{\rm int},K}\int_{0}^{\varepsilon t\wedge t_{0}}s^{4}Z_{s}^{(\infty)}(Z_{s}^{(\infty)}-1){\rm d}s=\hat{\bm{\sigma}}^{2}_{{\rm int},K}\int_{0}^{t\wedge(t_{0}/\varepsilon)}s^{2}(-\mathbf{m}_{\rm int}\varepsilon s)^{2}Z_{\varepsilon s}^{(\infty)}(Z_{\varepsilon s}^{(\infty)}-1){\rm d}s.$ (91)
We verify [10, assumptions (b) Theorem 1.4 in Chapter 7.1] with
$c_{1,1}(t)=\hat{\bm{\sigma}}^{2}_{{\rm int},k}\int_{0}^{t}s^{2}{\rm d}s\qquad\textrm{and}\qquad A_{n}^{11}(t)=\langle D_{k,\cdot}^{(\varepsilon)}\rangle_{t},\qquad\textrm{for}\qquad t\geq 0.$
Since $(\langle D_{k,\cdot}^{(\varepsilon)}\rangle_{t})_{t\geq 0}$ is a continuous process, [10, (1.16) in Theorem 1.4 in Chapter 7.1] is satisfied. Moreover, Theorem 3, Corollay 2 (i) and (91) imply that, for each fixed $t\geq 0$,
$\langle D_{k,\cdot}^{(\varepsilon)}\rangle_{t}\quad\textrm{converges to }\quad\hat{\bm{\sigma}}^{2}_{{\rm int},k}\int_{0}^{t}s^{2}{\rm d}s,\qquad\textrm{as}\qquad\varepsilon\downarrow 0,$
in probability (this is, [10, (1.19) in Theorem 1.4 in Chapter 7.1]). Finally, from the definition of $(D_{k,t}^{(\varepsilon)})_{t\geq 0}$ and by using (89), it follows that its jumps on $[0,T]$, for $0\leq T<\infty$, are uniformly bounded by $-\mathbf{m}_{\rm int}\varepsilon^{1/2}T^{2}k$, which implies [10, (1.17) in Theorem 1.4 in Chapter 7.1]. This concludes our proof. ∎
@@
$Y_{0}=0\qquad\text{and}\qquad Y_{t}:=t^{-1}\int_{0}^{t}s{\rm d}M_{s}^{\rm int},\quad\text{for}\quad t>0.$
Clearly, $(Y_{t})_{t\geq 0}$ is well-defined. For $\varepsilon>0$, define the stochastic process $(Y_{t}^{(\varepsilon)})_{t\geq 0}$ by letting $Y_{t}^{(\varepsilon)}=\varepsilon^{-\frac{1}{2}}Y_{\varepsilon t}$, for $t\geq 0$. Recall the definition of the stochastic process $(X_{t}^{(\varepsilon)})_{t\geq 0}$ in (10).
Lemma 21.
@@
$Y_{t}=-\int_{0}^{t}s^{-1}Y_{s}{\rm d}s+M_{t}^{\rm int},\quad\text{for}\quad t\geq 0.$ (92)
@@
$\mathbb{E}\Big{[}\sup_{s\leq t}Y_{s}^{2}\Big{]}\leq C(t\vee t^{2}),\quad\text{for}\quad t\leq T,$ (93)
@@
$\lim_{\varepsilon\downarrow 0}\mathbb{E}\Big{[}\sup_{t\leq T}\Big{|}X_{t}^{(\varepsilon)}-Y_{t}^{(\varepsilon)}\Big{|}\Big{]}=0.$ (94)
@@
By Corollary 2 (i), the predictable quadratic variation $(\langle D\rangle_{t})_{t\geq 0}$ of $(D_{t})_{t\geq 0}$ satisfies
$\mathbb{E}[\langle D\rangle_{t}]$ $=(\mathbf{m}_{\rm int})^{2}\bm{\sigma}^{2}_{{\rm int}}\mathbb{E}_{\infty}\Big{[}\int_{0}^{t\wedge t_{0}}s^{4}Z_{s}^{(\infty)}(Z_{s}^{(\infty)}-1){\rm d}s\Big{]}\leq C_{1}\int_{0}^{t}s^{2}{\rm d}s=\frac{C_{1}}{3}t^{3},$
@@
$\mathbb{E}[Y_{t}^{2}]=t^{-2}\mathbb{E}[D_{t}^{2}]=t^{-2}\mathbb{E}[\langle D\rangle_{t}]\leq\frac{C_{1}}{3}t.$ (95)
@@
The identity (92) follows by integration by parts (note that $t\mapsto 1/t$ is continuous and of finite variation on any closed interval $[a,b]$, for $0<a<b$). The only subtle point is the lack of regularity of $t\mapsto 1/t$ at 00. This technicality is overcome, by writing first the formula for $Y_{t}-Y_{r}$, for any $0<r\leq t$, i.e.,
$Y_{t}-Y_{r}=-\int_{r}^{t}s^{-2}\int_{0}^{s}u{\rm d}M_{u}^{\rm int}{\rm d}s+\int_{(r,t]}{\rm d}M_{s}^{\rm int}=-\int_{r}^{t}s^{-1}Y_{s}{\rm d}s+M_{t}^{\rm int}-M_{r}^{\rm int},$ (96)
@@
and then letting $r\downarrow 0$. Indeed, by using (95), we see that there exists a constant $0<C_{2}<\infty$ such that $\mathbb{E}[|\int_{0}^{t}s^{-1}Y_{s}{\rm ds}|]\leq C_{2}t^{1/2}$ and thus, $\int_{0}^{t}s^{-1}Y_{s}{\rm ds}$ is a well-defined Lebesgue integral for all $t\geq 0$ simultaneously, almost surely.
The estimate (93) follows from (92), [6, Lemma 10] and (83) in Lemma 16. Observe that by (88) and (92) we have that
$X_{t}^{(\varepsilon)}-Y_{t}^{(\varepsilon)}=-\varepsilon^{-1/2}\int_{0}^{\varepsilon t}s^{-1}(X_{s}-Y_{s}){\rm d}s+\varepsilon^{-1/2}M_{\varepsilon t}^{\rm br}+\varepsilon^{-1/2}U_{\varepsilon t},\quad\text{for}\quad 0\leq t\leq t^{\ast}/\varepsilon.$
@@
By [6, Lemma 10], $\sup_{s\leq t}|X_{s}^{(\varepsilon)}-Y_{s}^{(\varepsilon)}|\leq 2\varepsilon^{-1/2}\sup_{s\leq t}|M_{\varepsilon s}^{\rm br}+U_{\varepsilon s}|$. Then, by (82) and (86), there exists a constant $0<C<\infty$ such that $\mathbb{E}[\sup_{s\leq T}|X_{s}^{(\varepsilon)}-Y_{s}^{(\varepsilon)}|]\leq C(\varepsilon^{1/2}T\vee\varepsilon^{3/2}T^{2})$, for $0\leq T<\infty$ and $\varepsilon>0$ such that $T\leq t_{0}/\varepsilon$, which implies (94). This completes the proof. ∎
Finally, we prove Theorem 4.
Proof of Theorem 4.
@@
By (94) and the symmetry of the law of the standard Brownian motion $(W_{t})_{t\geq 0}$ it suffices to show that the process $(-Y^{(\varepsilon)}_{t})_{t\geq 0}$, converges weakly, in $\mathbb{D}([0,\infty),\mathbb{R})$, as $\varepsilon\downarrow 0$, to the process $(X_{t})_{t\geq 0}$. To do that, we follow the argument explained in the last part of the [29, proof of Theorem 1.1], which in turn is based in Steps 2-4 of the [30, proof of Lemma 4.8]. For $k>0$, define the stochastic process $(Y^{(\varepsilon)}_{k,t})_{t\geq 0}$ by letting
$Y^{(\varepsilon)}_{k,t}:=-(k^{-1}\mathbf{1}_{\{t\in[0,k]\}}+t^{-1}\mathbf{1}_{\{t\in(k,\infty)\}})D_{t}^{(\varepsilon)},\quad\text{for}\quad t\geq 0.$
@@
Recall that if $f:\mathbb{R}_{+}\rightarrow\mathbb{R}$ is continuous, then the mapping $w\mapsto f(w)$ is continuous from $\mathbb{D}([0,\infty),\mathbb{R})$ into itself. Hence, Proposition 3 implies that for any $k>0$, as $\varepsilon\downarrow 0$, the process $(Y^{(\varepsilon)}_{k,t})_{t\geq 0}$ converges weakly, in $\mathbb{D}([0,\infty),\mathbb{R})$, to the process $(X_{t}^{(k)})_{t\geq 0}$ defined by
$X^{(k)}_{t}:=-(tk^{-1}\mathbf{1}_{\{t\in[0,k]\}}+\mathbf{1}_{\{t\in(k,\infty)\}})X_{t},\quad\text{for}\quad t\geq 0.$
@@
Note that $Y^{(\varepsilon)}_{t}=-t^{-1}D_{t}^{(\varepsilon)}$ and thus, $Y^{(\varepsilon)}_{t}-Y^{(\varepsilon)}_{k,t}=(k^{-1}-t^{-1})\mathbf{1}_{\{t\in[0,k]\}}D_{t}^{(\varepsilon)}$, for $t\geq 0$. By (93), there exists a constant $0<C<\infty$ (independent of $\varepsilon$) such that
$\mathbb{E}\Big{[}\sup_{t\geq 0}|Y^{(\varepsilon)}_{t}-Y^{(\varepsilon)}_{k,t}|^{2}\Big{]}\leq 4\varepsilon\mathbb{E}\Big{[}\sup_{t\leq k}|Y^{(\varepsilon)}_{t}|^{2}\Big{]}\leq C(\varepsilon k\vee\varepsilon k^{2}).$ (97)
@@
On the other hand, $-X_{t}-X_{t}^{(k)}=(tk^{-1}-1)\mathbf{1}_{\{t\in[0,k]\}}X_{t}$, for $t\geq 0$. By Doob’s maximal inequality, there exists a constant $0<C<\infty$ (independent of $\varepsilon$) such that
$\mathbb{E}\Big{[}\sup_{t\geq 0}|-X_{t}-X_{t}^{(k)}|^{2}\Big{]}\leq 4\mathbb{E}\Big{[}\sup_{t\leq k}|X_{t}|^{2}\Big{]}\leq Ck.$ (98)
@@
Let $d_{\infty}^{\rm Sk}$ denote the Skorokhod metric on $\mathbb{D}(\mathbb{R}_{+},\mathbb{R})$ (here, $\mathbb{R}$ is equipped with the Euclidean distance) as defined in [19, Chapter VI, Section 1]. In particular, note that $d_{\infty}^{\rm Sk}(f,g)\leq\sup_{t\geq 0}|f(t)-g(t)|$, for any $f,g\in\mathbb{D}(\mathbb{R}_{+},\mathbb{R})$. Finally, it is straightforward to show that $(-Y^{(\varepsilon)}_{t})_{t\geq 0}$ converges to $(X_{t})_{t\geq 0}$. By [8, Theorem 2.1 in Chapter 1], it suffices to show that
$\lim_{\varepsilon\downarrow 0}|\mathbb{E}[F(-Y^{(\varepsilon)}_{t})_{t\geq 0})]-\mathbb{E}[F((X_{t})_{t\geq 0})]|=0,$
@@
for all all bounded, uniformly continuous function $F:\mathbb{D}(\mathbb{R}_{+},\mathbb{R})\rightarrow\mathbb{D}(\mathbb{R}_{+},\mathbb{R})$. The above follows from the convergence of $(Y^{(\varepsilon)}_{k,t})_{t\geq 0}$ towards $(X_{t}^{(k)})_{t\geq 0}$ (i.e., for any $k>0$, $|\mathbb{E}[F((Y^{(\varepsilon)}_{k,t})_{t\geq 0})]-\mathbb{E}[F((X_{t}^{(k)})_{t\geq 0})]|\rightarrow 0$, as $\varepsilon\downarrow 0$), (97), (98), the triangle inequality, the uniform continuity of $F$ and the Markov inequality and the above discussion (the argument based on the addition and subtraction of intermediate terms). This concludes the proof. ∎

### v2

@@ 3.1 Integral representation of BPI processes
$\mathbb{E}\Big{[}\sup_{s\leq t}(D_{s}-D_{k,s})^{2}\Big{]}\leq C\bm{\sigma}^{2}_{{\rm int},k}t^{3},\quad\text{for}\quad t\leq T,$ (3.106)
@@
$\langle D_{k,\cdot}-D_{\cdot}\rangle_{t}=(\mathbf{m}_{\rm int})^{2}\bm{\sigma}^{2}_{{\rm int},k}\int_{0}^{t\wedge t^{\ast}}s^{4}Z_{s}^{(\infty)}\left(Z_{s}^{(\infty)}-1\right){\rm d}s,\quad\text{for}\quad t\geq 0.$ (3.107)
Then, for any $0\leq T<\infty$, Corollary 3.11 (i) implies that there exists a constant $0<C<\infty$ such that
$\mathbb{E}[\langle D_{k,\cdot}-D_{\cdot}\rangle_{t}]$ $\leq C\bm{\sigma}^{2}_{{\rm int},k}t^{3},\quad\text{for}\quad t\leq T.$ (3.108)
@@
$D_{t}^{(\varepsilon)}:=-\varepsilon^{-\frac{3}{2}}D_{\varepsilon t},\qquad\textrm{ for }\quad t\geq 0,$ (3.109)
@@
$\widehat{X}_{0}=0\qquad\text{and}\qquad\widehat{X}_{t}=\bm{\sigma}_{{\rm int}}\int_{0}^{t}u{\rm d}W_{u},\qquad\text{for}\quad t>0,$ (3.110)
@@
$\widehat{X}_{0}^{(k)}=0\qquad\text{and}\qquad\widehat{X}_{t}^{(k)}=\hat{\bm{\sigma}}_{{\rm int},k}\int_{0}^{t}u{\rm d}W_{u},\qquad\text{for}\quad t>0,$ (3.111)
where $\hat{\bm{\sigma}}^{2}_{{\rm int},k}\coloneqq c+\sum_{i=1}^{\lfloor k\rfloor}i^{2}b_{i}$. Recall that convergence in $\mathbb{D}([0,\infty),\mathbb{R})$ to a continuous limit is equivalent to uniform convergence on compact subsets of $0,\infty)$, see for instance Proposition 1.17 in Chapter VI of [18]. Therefore, since by ($\bf H4$}), $\hat{\bm{\sigma}}^{2}_{{\rm int},k}\rightarrow\bm{\sigma}^{2}_{{\rm int}}$, as $k\rightarrow\infty$, it is not difficult to see that Lemma 3.18 and (3.111) imply the result in Proposition 3.19.
To prove that $(D_{k,t}^{(\varepsilon)})_{t\geq 0}$ converges weakly towards $(\widehat{X}_{t}^{(k)})_{t\geq 0}$, we use Theorem 1.4 in Chapter 7.1 of [10]. Note that $(D_{k,t}^{(\varepsilon)})_{t\geq 0}$ is a well-defined locally square-integrable martingale. In particular, by (3.105), its predictable quadratic variation $(\langle D_{k,\cdot}^{(\varepsilon)}\rangle_{t})_{t\geq 0}$ is given by
$\langle D_{k,\cdot}^{(\varepsilon)}\rangle_{t}$ $=\varepsilon^{-3}(\mathbf{m}_{\rm int})^{2}\hat{\bm{\sigma}}^{2}_{{\rm int},K}\int_{0}^{\varepsilon t\wedge t_{0}}s^{4}Z_{s}^{(\infty)}(Z_{s}^{(\infty)}-1){\rm d}s$
$=\hat{\bm{\sigma}}^{2}_{{\rm int},K}\int_{0}^{t\wedge(t_{0}/\varepsilon)}s^{2}(-\mathbf{m}_{\rm int}\varepsilon s)^{2}Z_{\varepsilon s}^{(\infty)}(Z_{\varepsilon s}^{(\infty)}-1){\rm d}s.$ (3.112)
We verify assumptions (b) Theorem 1.4 in Chapter 7.1 of [10] with
$c_{1,1}(t)=\hat{\bm{\sigma}}^{2}_{{\rm int},k}\int_{0}^{t}s^{2}{\rm d}s\qquad\textrm{and}\qquad A_{n}^{11}(t)=\langle D_{k,\cdot}^{(\varepsilon)}\rangle_{t},\qquad\textrm{for}\qquad t\geq 0.$ (3.113)
Since $(\langle D_{k,\cdot}^{(\varepsilon)}\rangle_{t})_{t\geq 0}$ is a continuous process, (1.16) in Theorem 1.4 in Chapter 7.1 of [10] is satisfied. Moreover, Theorem 2.2, Corollay 3.11 (i) and (3.4) imply that, for each fixed $t\geq 0$,
$\langle D_{k,\cdot}^{(\varepsilon)}\rangle_{t}\quad\textrm{converges to }\quad\hat{\bm{\sigma}}^{2}_{{\rm int},k}\int_{0}^{t}s^{2}{\rm d}s,\qquad\textrm{as}\qquad\varepsilon\downarrow 0,$ (3.114)
in probability (this is, (1.19) in Theorem 1.4 in Chapter 7.1 of [10]). Finally, from the definition of $(D_{k,t}^{(\varepsilon)})_{t\geq 0}$ and by using (3.105), it follows that its jumps on $[0,T]$, for $0\leq T<\infty$, are uniformly bounded by $\mathbf{m}_{\rm int}\varepsilon^{1/2}T^{2}k$, which implies (1.17) in Theorem 1.4 in Chapter 7.1 of [10]. This concludes our proof. ∎
@@
$Y_{0}=0\qquad\text{and}\qquad Y_{t}:=t^{-1}\int_{0}^{t}s{\rm d}M_{s}^{\rm int},\quad\text{for}\quad t>0.$ (3.115)
Clearly, $(Y_{t})_{t\geq 0}$ is well-defined. For $\varepsilon>0$, define the stochastic process $(Y_{t}^{(\varepsilon)})_{t\geq 0}$ by letting $Y_{t}^{(\varepsilon)}=\varepsilon^{-\frac{1}{2}}Y_{\varepsilon t}$, for $t\geq 0$. Recall the definition of the stochastic process $(X_{t}^{(\varepsilon)})_{t\geq 0}$ in (2.5).
Lemma 3.20.
@@
$Y_{t}=-\int_{0}^{t}s^{-1}Y_{s}{\rm d}s+M_{t}^{\rm int},\quad\text{for}\quad t\geq 0.$ (3.116)
@@
$\mathbb{E}\Big{[}\sup_{s\leq t}Y_{s}^{2}\Big{]}\leq C(t\vee t^{2}),\quad\text{for}\quad t\leq T,$ (3.117)
@@
$\lim_{\varepsilon\downarrow 0}\mathbb{E}\Big{[}\sup_{t\leq T}\Big{|}X_{t}^{(\varepsilon)}-Y_{t}^{(\varepsilon)}\Big{|}\Big{]}=0.$ (3.118)
@@
By Corollary 3.11 (i), the predictable quadratic variation $(\langle D\rangle_{t})_{t\geq 0}$ of $(D_{t})_{t\geq 0}$ satisfies
$\mathbb{E}[\langle D\rangle_{t}]$ $=(\mathbf{m}_{\rm int})^{2}\bm{\sigma}^{2}_{{\rm int}}\mathbb{E}_{\infty}\Big{[}\int_{0}^{t\wedge t^{\ast}}s^{4}Z_{s}^{(\infty)}(Z_{s}^{(\infty)}-1){\rm d}s\Big{]}\leq C_{1}\int_{0}^{t}s^{2}{\rm d}s=\frac{C_{1}}{3}t^{3},$ (3.119)
@@
$\mathbb{E}[Y_{t}^{2}]=t^{-2}\mathbb{E}[D_{t}^{2}]=t^{-2}\mathbb{E}[\langle D\rangle_{t}]\leq\frac{C_{1}}{3}t.$ (3.120)
@@
The identity (3.116) follows by integration by parts (note that $t\mapsto 1/t$ is continuous and of finite variation on any closed interval $[a,b]$, for $0<a<b$). The only subtle point is the lack of regularity of $t\mapsto 1/t$ at 0. This technicality is overcome, by writing first the formula for $Y_{t}-Y_{r}$, for any $0<r\leq t$, i.e.,
$Y_{t}-Y_{r}=-\int_{r}^{t}s^{-2}\int_{0}^{s}u{\rm d}M_{u}^{\rm int}{\rm d}s+\int_{(r,t]}{\rm d}M_{s}^{\rm int}=-\int_{r}^{t}s^{-1}Y_{s}{\rm d}s+M_{t}^{\rm int}-M_{r}^{\rm int},$ (3.121)
@@
and then letting $r\downarrow 0$. Indeed, by using (3.120), we see that there exists a constant $0<C_{2}<\infty$ such that $\mathbb{E}[|\int_{0}^{t}s^{-1}Y_{s}{\rm ds}|]\leq C_{2}t^{1/2}$ and thus, $\int_{0}^{t}s^{-1}Y_{s}{\rm ds}$ is a well-defined Lebesgue integral for all $t\geq 0$ simultaneously, almost surely.
The estimate (3.117) follows from (3.116), Lemma 10 in [6] and (3.84) in Lemma 3.14. Observe that by (3.99) and (3.116) we have that
$X_{t}^{(\varepsilon)}-Y_{t}^{(\varepsilon)}=-\varepsilon^{-1/2}\int_{0}^{\varepsilon t}s^{-1}(X_{s}-Y_{s}){\rm d}s+\varepsilon^{-1/2}M_{\varepsilon t}^{\rm br}+\varepsilon^{-1/2}U_{\varepsilon t},\quad\text{for}\quad 0\leq t\leq t^{\ast}/\varepsilon.$ (3.122)
@@
By Lemma 10 in [6], $\sup_{s\leq t}|X_{s}^{(\varepsilon)}-Y_{s}^{(\varepsilon)}|\leq 2\varepsilon^{-1/2}\sup_{s\leq t}|M_{\varepsilon s}^{\rm br}+U_{\varepsilon s}|$. Then, by (3.83) and (3.96), there exists a constant $0<C<\infty$ such that $\mathbb{E}[\sup_{s\leq T}|X_{s}^{(\varepsilon)}-Y_{s}^{(\varepsilon)}|]\leq C(\varepsilon^{1/2}T\vee\varepsilon^{3/2}T^{2})$, for $0\leq T<\infty$ and $\varepsilon>0$ such that $T\leq t^{\ast}/\varepsilon$, which implies (3.118). This completes the proof. ∎
Finally, we prove Theorem 2.3.
Proof of Theorem 2.3.
@@
By (3.118) and the symmetry of the law of the standard Brownian motion $(W_{t})_{t\geq 0}$ it suffices to show that the process $(-Y^{(\varepsilon)}_{t})_{t\geq 0}$, converges weakly, in $\mathbb{D}([0,\infty),\mathbb{R})$, as $\varepsilon\downarrow 0$, to the process $(X_{t})_{t\geq 0}$. To do that, we follow the argument explained in the last part of the proof of Theorem 1.1 of [28], which in turn is based in Steps 2-4 of the proof of Lemma 4.8 of [29]. For $k>0$, define the stochastic process $(Y^{(\varepsilon)}_{k,t})_{t\geq 0}$ by letting
$Y^{(\varepsilon)}_{k,t}:=-(k^{-1}\mathbf{1}_{\{t\in[0,k]\}}+t^{-1}\mathbf{1}_{\{t\in(k,\infty)\}})D_{t}^{(\varepsilon)},\quad\text{for}\quad t\geq 0.$ (3.123)
@@
Recall that if $f:\mathbb{R}_{+}\rightarrow\mathbb{R}$ is continuous, then the mapping $w\mapsto f(w)$ is continuous from $\mathbb{D}([0,\infty),\mathbb{R})$ into itself. Hence, Proposition 3.19 implies that for any $k>0$, as $\varepsilon\downarrow 0$, the process $(Y^{(\varepsilon)}_{k,t})_{t\geq 0}$ converges weakly, in $\mathbb{D}([0,\infty),\mathbb{R})$, to the process $(X_{t}^{(k)})_{t\geq 0}$ defined by
$X^{(k)}_{t}:=-(tk^{-1}\mathbf{1}_{\{t\in[0,k]\}}+\mathbf{1}_{\{t\in(k,\infty)\}})X_{t},\quad\text{for}\quad t\geq 0.$ (3.124)
@@
Note that $Y^{(\varepsilon)}_{t}=-t^{-1}D_{t}^{(\varepsilon)}$ and thus, $Y^{(\varepsilon)}_{t}-Y^{(\varepsilon)}_{k,t}=(k^{-1}-t^{-1})\mathbf{1}_{\{t\in[0,k]\}}D_{t}^{(\varepsilon)}$, for $t\geq 0$. By (3.117), there exists a constant $0<C<\infty$ (independent of $\varepsilon$) such that
$\mathbb{E}\Big{[}\sup_{t\geq 0}|Y^{(\varepsilon)}_{t}-Y^{(\varepsilon)}_{k,t}|^{2}\Big{]}\leq 4\varepsilon\mathbb{E}\Big{[}\sup_{t\leq k}|Y^{(\varepsilon)}_{t}|^{2}\Big{]}\leq C(\varepsilon k\vee\varepsilon k^{2}).$ (3.125)
@@
On the other hand, $X_{t}-X_{t}^{(k)}=(tk^{-1}-1)\mathbf{1}_{\{t\in[0,k]\}}X_{t}$, for $t\geq 0$. By Doob’s maximal inequality, there exists a constant $0<C<\infty$ (independent of $\varepsilon$) such that
$\mathbb{E}\Big{[}\sup_{t\geq 0}|-X_{t}-X_{t}^{(k)}|^{2}\Big{]}\leq 4\mathbb{E}\Big{[}\sup_{t\leq k}|X_{t}|^{2}\Big{]}\leq Ck.$ (3.126)
@@
Let $d_{\infty}^{\rm Sk}$ denote the Skorokhod metric on $\mathbb{D}(\mathbb{R}_{+},\mathbb{R})$ (here, $\mathbb{R}$ is equipped with the Euclidean distance) as defined in Section 1 of Chapter VI in [18]. In particular, note that $d_{\infty}^{\rm Sk}(f,g)\leq\sup_{t\geq 0}|f(t)-g(t)|$, for any $f,g\in\mathbb{D}(\mathbb{R}_{+},\mathbb{R})$. Finally, it is straightforward to show that $(-Y^{(\varepsilon)}_{t})_{t\geq 0}$ converges to $(X_{t})_{t\geq 0}$. By Theorem 2.1 in Chapter 1 of [8], it suffices to show that
$\lim_{\varepsilon\downarrow 0}|\mathbb{E}[F(-Y^{(\varepsilon)}_{t})_{t\geq 0})]-\mathbb{E}[F((X_{t})_{t\geq 0})]|=0,$ (3.127)
@@
for all all bounded, uniformly continuous function $F:\mathbb{D}(\mathbb{R}_{+},\mathbb{R})\rightarrow\mathbb{D}(\mathbb{R}_{+},\mathbb{R})$. The above follows from the convergence of $(Y^{(\varepsilon)}_{k,t})_{t\geq 0}$ towards $(X_{t}^{(k)})_{t\geq 0}$ (i.e., for any $k>0$, $|\mathbb{E}[F((Y^{(\varepsilon)}_{k,t})_{t\geq 0})]-\mathbb{E}[F((X_{t}^{(k)})_{t\geq 0})]|\rightarrow 0$, as $\varepsilon\downarrow 0$), (3.125), (3.126), the triangle inequality, the uniform continuity of $F$, the Markov inequality and the above discussion (the argument is based on the addition and subtraction of intermediate terms). This concludes the proof. ∎